Lane Company manufactures a single product that requires a great deal of hand labor. Overhead cost is applied on the basis ofReq 1 Reg 2 Req ЗА Req 3B Req 4 Compute the predetermined overhead rate for the year. Break the rate down into variable and fReq 1 Reg 2 Req 3A Req 3B Reg 4 Prepare a standard cost card for the companys product. (Round your answers to 2 decimal placReq 1 Reg 2 Req ЗА Req 3B Req 4 Compute the standard direct labor-hours allowed for the years production. Standard direct laReq 1 Req 2 Req ЗА Req 3B Req 4 Complete the following Manufacturing Overhead T-account for the year. Manufacturing OverheadReg 1 Reg 2 Req ЗА Req 3B Req 4 Determine the reason for the underapplied or overapplied overhead from (3) above by computing
Answer #1

Solution 1 and 2:

G H 5 Solution 1: | Computation of Predetermined overhead rate - Lane Company Particulars Amount Budgeted variable manufactur

Solution 3a:

Standard hours allowed = 36000*1.5 = 54,000 hours

Solution 3b:

R solution 3b: Particulars Actual costs incurred(87750+321750) Manufacturing Overhead Debit Particulars Credit $409,500 $453,

Solution 4:

Variable overhead actual rate = $87750/ 58500 = $1.50 per hour

Variable overhead rate variance = (SP - AP) *Actual Hours = ($2.20 - $1.50) * 58500 = $40,950 Favorable

Variable overhead Efficiency variance = (Standard hours - Actual Hours) *SP = (54000 - 58500)*$2.20 = $9,900 (Unfavorable)

Fixed overhead budget Variance = Budgeted Fixed Overhead - Actual Fixed overhead = $279000 - $321750 = $42,750 (unfavorable)

Fixed Overhead Volume Variance = ($6.20*54000) - $279000 = 55,800 Favorable